OpenStudy (anonymous):

A quadratic function passes through the point (-11, -21) and has a vertex at (-8, -48). Write the function in the form y = a(x - p)2 + q. What are the values of a, p and q? The value of 'a' is

OpenStudy (saifoo.khan):

This is same as, a(x-h)^2 + k

OpenStudy (saifoo.khan):

a(x+11)^2 - 21

OpenStudy (anonymous):

i know p and Q but i dont remeber what do u put if there is not number for a just 1?

OpenStudy (anonymous):

and also would p be -8 since it says the vertex is -8,-48 ?

OpenStudy (saifoo.khan):

i made a mistake! :( kill me..

OpenStudy (saifoo.khan):

a(x+8)^2 - 48 It will be this!

OpenStudy (anonymous):

hahahah i wont just wondering , cuz im not so good at it as u seee lol , :) thank you soo much so A can be considered 1?

OpenStudy (saifoo.khan):

Lol, you are way better than others.. Now after you have a(x+8)^2 - 48 use it to find the values for a. \[ y = a(x+8)^2 - 48\]\[-21 = a((-11)+8)^2 - 48\]

OpenStudy (anonymous):

you seem to be extremly good in math helped me alot , thank you : )

OpenStudy (saifoo.khan):

Thankyou. Math is fairly easy. ;) Dont forget to hit good answer. Tx. @irena95

OpenStudy (anonymous):

best answer ;p

OpenStudy (saifoo.khan):

Haha. thanks.

OpenStudy (saifoo.khan):

I used to be "good answer" about 10 days ago..